OpenStudy (anonymous):

Use the formula for future value, A = P(1 + rt), to find the missing quantity. P = $1,800; r = 3%; t = 5 years

OpenStudy (mr.math):

What's the missing quantity here?

OpenStudy (anonymous):

I got 1,565.21. Because I was missing A. Is that correct? I keep doing these problems wrong. I check them on here.

OpenStudy (mr.math):

Shouldn't it be A=P(1+r)^t?

OpenStudy (anonymous):

I don't know? The question had the formula I am supposed to use.

OpenStudy (anonymous):

Is my answer correct?

OpenStudy (mr.math):

Well if this is the formula, then \(A=1800(1+0.03\times 5)=$2070.\)

OpenStudy (anonymous):

Thank you, again! Your help is much appreciated!

OpenStudy (mr.math):

you're very welcome! :D

OpenStudy (anonymous):

I am not sure what I did that I got completely the wrong answer? Can you please show me your steps again? I am having the worst time with these problems. Do you mind?

OpenStudy (mr.math):

I don't mind, but actually there are no steps if you are allowed to use a calculator other than substituting into the given formula and then computing the value using the calculator. You need to be careful about the brackets though.

OpenStudy (anonymous):

Because I also have this problem and I think I got this wrong, too. The principal P is borrowed at simple interest rate r for a period of time t. Find the loan's future value, A, or the total amount due at time t. P = $2,000, r = 6%, t = 3 years (I got 2,360.)

OpenStudy (mr.math):

CORRECT!! :D

OpenStudy (anonymous):

REALLY?! Oh man. THANK YOU! You definitely are helping me!

OpenStudy (mr.math):

But I still don't know why they are using this formula, I don't want to confuse you anyway :D Best of Luck!

OpenStudy (anonymous):

Thank you! Chances are I will be putting on some more questions, so I'm becoming a fan. I hope you come back and help me!

OpenStudy (mr.math):

I will. Put them in separate posts!

OpenStudy (anonymous):

Thank you!